Can someone please help me with this?
Colin is buying dirt to fill a garden bed that is 9 feet by 16 feet. If he wants to fill it to depth of 4 in. how many cubic yards of dirt does he need?
Thank you!

Answer by nerdybill(7384)

Colin is buying dirt to fill a garden bed that is 9 feet by 16 feet. If he wants to fill it to depth of 4 in. how many cubic yards of dirt does he need?
.
1 cubic yard = 3*3*3 = 27 cubic feet
4 in = 1/3 feet
.
volume = 1/3*9*16 = 3*16 = 48 cubic feet
volume = 48/27 = 1.78 cubic yards